Alex2007:

If f(x)=3x+2 and g(x)=x2−x, find the value. g(−3)=

worthy13:

f(2)+1

carlosbigbain:

To find the value of g(-3), we can substitute -3 into the function g(x) and simplify. Let's do it step by step: First, let's write out the function g(x) as g(x) = x^2 - x. Now, substitute -3 for x in the function: g(-3) = (-3)^2 - (-3). Simplifying further: g(-3) = 9 + 3. And the final answer is: g(-3) = 12. So, the value of g(-3) is 12.

Tsukage:

The answer is a number :)

Alex2007:

@carlosbigbain wrote:
To find the value of g(-3), we can substitute -3 into the function g(x) and simplify. Let's do it step by step: First, let's write out the function g(x) as g(x) = x^2 - x. Now, substitute -3 for x in the function: g(-3) = (-3)^2 - (-3). Simplifying further: g(-3) = 9 + 3. And the final answer is: g(-3) = 12. So, the value of g(-3) is 12.
thank you

carlosbigbain:

np medals is appreciated

Tsukage:

Sobbing